OpenStudy (jeremiah):

Solve x2 + 18x - 40 = 0.

OpenStudy (cwrw238):

first factor the trinomial it will be of the form (x - a)(x + b) where a and b are constants to find a and b: you need two numbers whose product is -40 and whose sum is + 18

OpenStudy (cwrw238):

have you had experience of factoring these?

OpenStudy (jeremiah):

no I don't get it at all

OpenStudy (cwrw238):

ok the 2 numbers are + 20 and -2 20*-2 = -40 and 20 - 2 = 18 so we have (x + 20)( x - 2) = 0 when 2 items are multiplied together to give zero either one of them must equal zero so either x + 20 = 0 or x - 2 = 0 that gives x = -20 or x = 2 solution is -20, 2

OpenStudy (cwrw238):

hope this helps

OpenStudy (jeremiah):

thanks alot

OpenStudy (cwrw238):

yw